# Closed Orbits of Dynamically Convex Reeb Flows: Towards the HZ- and Multiplicity Conjectures

### v1

@@ 1.1. Introduction
Recall that by the celebrated theorem of Franks a Hamiltonian diffeomorphism of $S^{2}$ with more than two fixed points must have infinitely many periodic orbits; see [Fr92, Fr96, LeC]. A generalization of Franks’ theorem to higher dimensions was conjectured by Hofer and Zehnder in [HZ, p. 263], asserting that a Hamiltonian diffeomorphism of ${\mathbb{C}}{\mathbb{P}}^{n}$ with more than $n+1$ fixed points must have infinitely many periodic orbits. (Hence, the name “HZ-conjecture”.) Here, by Arnold’s conjecture, $n+1$ is the minimal possible number of fixed points, [Fl89b, Fo, FW]. For instance, $n+1=2$ for ${\mathbb{C}}{\mathbb{P}}^{1}=S^{2}$. The HZ-conjecture was proved in [Sh22] under a minor additional non-degeneracy type condition but for a broader class of manifolds than complex projective spaces; see also, e.g., [Al22a, Al22b, Al23, ABF, AL, Ba15, Ba17, Ba18, ÇGG22, GG14, GG16, Gü13, Su21] for other relevant results, conjectures and more context.
@@
From a slightly different perspective, the $n$-or-$\infty$ conjecture is a question about Reeb pseudo-rotations. In the context of symplectic dynamics, the two competing definitions of a pseudo-rotation are that this is a Hamiltonian system (a Hamiltonian diffeomorphism or a Reeb flow) with finitely many periodic orbits or, alternatively, with the minimal possible number of periodic orbits (perhaps, counted with multiplicity); see [GG18, Sh22, ÇG2M23]. Hypothetically, the two definitions are equivalent and Theorem B establishes this fact for a certain class of Reeb flows, just as [Sh22] did for Hamiltonian diffeomorphisms of ${\mathbb{C}}{\mathbb{P}}^{n}$ and some other manifolds. (Interestingly, this is no longer true for symplectomorphisms, [ABF].) Overall, perhaps expectedly so, there is a considerable similarity between the properties of Hamiltonian and Reeb pseudo-rotations; see, e.g., [AK, LRS] vs. [Ka] or [GG18] vs. [ÇG2M23].
@@
Finally, the case of dimension $32n-1=$ 3 deserves a separate discussion. According to the so called 2-or-infinity conjecture, a Reeb flow on a closed 3-manifold has either exactly two or infinitely many prime closed orbits. This conjecture is almost proved in general, and lens spaces are the only manifolds admitting Reeb pseudo-rotations; see [CGHu, CGHP, CGH2L, CDR]. Furthermore, the conjecture has been completely proved for $S^{3}$. (See also [GH2M] for a different proof for $S^{3}$.) Hence, both (C-M) and (C-HZ) parts of our $n$-or-$\infty$ conjecture hold in this case. However, one cannot expect this clear-cut dichotomy to hold in higher dimensions, even for tight or fillable contact manifolds. Indeed, while it is reasonable to expect that a majority of such contact manifolds do not admit Reeb pseudo-rotations in all dimensions just as in dimension three, it is not clear how to rigorously state this conjecture and a description of manifolds that do seems beyond our reach. Moreover, the lower bound on the number of orbits clearly depends on the manifold when $n>2$.
@@
We use Theorem C to show that every homologically visible (over ${\mathbb{Q}}$) closed orbit has one-dimensional equivariant local symplectic homology. This allows to assign such a closed orbit a degree and treat the orbit as if it were non-degenerate. We show that the degrees of such orbits form the increasing sequence $n+1,n+3,\ldots$ and the degrees of bars are $n,n+2,\ldots$. All closed orbits have distinct period (action) and the same ratio of the action and the mean index. (See Corollary D and Theorem E.) In other words, the (non-equivariant) symplectic homology persistence module and the orbit degrees behave as they would for an irrational ellipsoid.
@@ 1.2. Main results
When $x$ is non-degenerate and prime, we call it non-alternating if the number, counted with algebraic multiplicity, of real negative eigenvalues in $(-1,0)$ of its Poincaré return map is even. Otherwise, we call $x$ alternating. A non-degenerate even iterate of an alternating closed orbit is called bad; all other non-degenerate closed orbits are called good. When $M$ is centrally symmetric, we call $x$ symmetric if $-x(t)=x(t+T/2)$, where $T$ is the period of $x$, i.e., $x$ and $-x$ are the same orbit up to the choice of initial condition.
@@
This theorem is proved in Section 4.1. The new point here is, of course, (M1), establishing Ekeland’s conjecture and improving previously known multiplicity lower bounds by roughly a factor of two. Part (M2), included only for the sake of completeness and expository reasons, readily follows from the standard proof of the multiplicity theorem for non-degenerate dynamically convex flows and does not really need (M1); cf. the proof of Corollary 3.7 in Section 3.2.
@@ 2.1. Conventions and notation
A non-degenerate closed orbit $x$ is said to be non-alternating if the number (counted with algebraic multiplicity) of eigenvalues in $(-1,0)$ of the return map is even. Otherwise, we call $x$ alternating. When $x$ is strongly non-degenerate, these conditions are equivalent to that all indices $\mu\big{(}x^{k}\big{)}$, $k\in{\mathbb{N}}$, have the same parity or that the parity alternates. (The notion of (non-)alternating closed orbit is not particularly useful when $x^{k}$ is allowed to be degenerate; see Example 2.13.) The parity of a closed non-degenerate orbit $x$ is simply the parity of $\mu(x)$. A non-degenerate closed orbit $x$ is said to be bad if it is an even iterate of an alternating prime orbit, and good otherwise.
@@
This condition is obviously satisfied in two cases we are interested in here: when $\alpha$ is non-degenerate and when the Reeb flow has only finitely many prime closed orbits. The role of this condition is rather expository than conceptual, and most of the discussion in next two sections carries over to the general case with only technical changes. For instance, the symplectic homology is still a persistence module (see Section 2.2) when ${\mathcal{S}}$ is only closed and nowhere dense and $V^{s}$ might be infinite-dimensional, once (PM2) in the definition of a persistence module is suitably modified; [ÇG2M24].
@@ 2.3. Filtered symplectic homology
The construction of symplectic homology can be carried over ${\mathbb{Z}}$ or any other ring, and hence we have the universal coefficient theorem. In particular, for a fixed $t$, in every degree $m$ we have
@@
Recall in this connection that $\operatorname{SH}(W;{\mathbb{F}})$ vanishes, for instance, whenever $W$ is displaceable in $\widehat{W}$. For example, $\operatorname{SH}(W)=0$ when $W$ is a star-shaped domain in $\widehat{W}={\mathbb{R}}^{2n}$. In this case, vanishing of $\operatorname{SH}(W)$ is established in [Vi99]. The general case is proved in [CFO] via Rabinowitz–Floer homology and direct proofs are given in [Su16] and [GS]. Strictly speaking, Theorem 2.3 is proved in [GS] for ${\mathbb{F}}={\mathbb{Q}}$, but the argument carries over word-for-word to any unital ring.
@@
It is worth keeping in mind that this description does not carry over to a field of positive characteristic $p\geq 2$. The reason is that $x^{p}$ can contribute infinitely many generators to the complex over ${\mathbb{F}}_{p}$. For instance, if $x^{p}$ is non-degenerate and the only orbit with action in $I$ and $x$ is simple, $\operatorname{CH}^{I}(W;{\mathbb{F}}_{p})$ is isomorphic to the homology of the infinite-dimensional lens space $S^{2\infty-1}/{\mathbb{Z}}_{p}$ over ${\mathbb{F}}_{p}$.
@@ 2.4.2. Behavior under iterations
The notion of an (non-)alternating closed orbit is not particularly useful in the setting where $x$ is not strongly non-degenerate and the return map $\varphi$ has roots of unity of even degree among its eigenvalues. The reason is that in this case the behavior of the local symplectic homology is not determined by the eigenvalues and depends on higher order terms. For instance, in the spirit of Example 2.12, let $R_{\pi}\colon{\mathbb{R}}^{2}\to{\mathbb{R}}^{2}$ be the counterclockwise rotation by $\pi$, i.e., central symmetry, viewed as an element of $\operatorname{\widetilde{Sp}}(2)$. Consider, the Hamiltonians $H_{0}=p^{4}-q^{4}$ and $H_{1}=(p^{2}+q^{2})^{2}$ on ${\mathbb{R}}^{2}$, where $(p,q)$ are the Darboux coordinates. The flows of $H_{0}$ and $H_{1}$ commute with $R_{\pi}$. Let $\varphi_{i}$, for $i=0,1$, be obtained by composing $R_{\pi}$ with the time-one flow of $H_{i}$ and $x_{i}$ be a closed Reeb orbit with the return map $\varphi_{i}$. Clearly, $\operatorname{CH}(x_{i};{\mathbb{Q}})=\operatorname{HF}(\varphi_{i})={\mathbb{Q}}$ is supported in degree one. However, one can show that $\operatorname{CH}\big{(}x_{0}^{2};{\mathbb{Q}}\big{)}=0$ and $\operatorname{HF}\big{(}\varphi_{0}^{2};{\mathbb{Q}}\big{)}={\mathbb{Q}}$ (supported in degree two). At the same time, $\operatorname{CH}\big{(}x_{1}^{2};{\mathbb{Q}}\big{)}=\operatorname{HF}\big{(}\varphi_{1}^{2};{\mathbb{Q}}\big{)}={\mathbb{Q}}$ (supported in degree three).
@@ 2.4.3. From local to filtered homology
where, when $0\in I$, we treat $W$ as a collection of “orbit” with action 0 as described above. This spectral sequence is simply associated with the action filtration; see, e.g., [GG20, Sec. 2].
@@ 2.5. The beginning and end maps
For a fixed field ${\mathbb{F}}$, denote by ${\mathcal{B}}$ or ${\mathcal{B}}({\mathbb{F}})$ the collection of bars of the graded persistence module $\operatorname{SH}(W;{\mathbb{F}})$. Assume that $\operatorname{SH}^{\infty}(W;{\mathbb{F}})=0$. Then every bar in ${\mathcal{B}}$ is finite by Theorem 2.3. Furthermore, let ${\mathcal{P}}={\mathcal{P}}({\mathbb{F}})$ be the set of ${\mathbb{F}}$-visible closed Reeb orbits. As in the previous section, let us also formally add to ${\mathcal{P}}$ the set of the “beginning” of the bars of the form $I=(0,a]$ for some $a>0$, denoting the resulting set by ${\mathcal{P}}^{\prime}$. These additional elements $x$ of ${\mathcal{P}}^{\prime}$ are simply in one-to-one correspondence with such bars. By definition, each of such “orbits” $x$ has symplectic homology equal to ${\mathbb{F}}$ and concentrated in $\deg(I)$, and zero action. (We also set $\mu_{-}(x)=\mu_{+}(x):=\deg(x)$.) When $W$ is star-shaped, we are adding just one extra element corresponding to the domain $W$, which we denote by $[W]$. This element has action zero and its supported in degree $n$.
@@ 3.1.3. Conley–Zehnder index: closed orbits
Next, let $x\colon S_{T}^{1}:={\mathbb{R}}/T{\mathbb{Z}}\to M$ be a closed orbit of the Reeb flow, which we do not require to be prime. The pull-back ${\mathfrak{s}}_{M}\circ x$ of ${\mathfrak{s}}_{M}$ by $x$ is a non-vanishing section of the pull-back $x^{*}L_{M}$ over $S_{T}^{1}$ and $x^{*}L_{M}$ is in turn the top complex wedge of $x^{*}\xi$. Pick any Kähler trivialization of $x^{*}\xi$ which, up to homotopy, gives rise to the trivialization ${\mathfrak{s}}_{M}\circ x$ of $x^{*}L$, i.e., such that its top complex wedge is ${\mathfrak{s}}_{M}\circ x$. (Such a trivialization is unique up to homotopy.) Then we use this trivialization to turn the linearized Reeb low $D\varphi^{t}$, $t\in[0,T]$, along $x$ into a path $\Phi(t)\in\operatorname{Sp}(2m)$ with $m=n-1$ starting at the identity, and hence an element $\Phi\in\operatorname{\widetilde{Sp}}(2m)$. By definition, $\mu_{\pm}(x):=\mu_{\pm}(\Phi)$ and $\operatorname{\hat{\mu}}(x):=\operatorname{\hat{\mu}}(\Phi)$. In particular, when $x$ is non-degenerate, $\mu(x):=\operatorname{\hat{\mu}}(\Phi)$. Furthermore, $\operatorname{\hat{\mu}}$ is homogenenous, i.e.,

### v2

@@ 1.1. Introduction
Recall that by the celebrated theorem of Franks a Hamiltonian diffeomorphism of $S^{2}$ with more than two fixed points must have infinitely many periodic orbits; see [Fr92, Fr96, LeC]. A generalization of Franks’ theorem to higher dimensions was conjectured by Hofer and Zehnder in [HZ, p. 263], asserting that a Hamiltonian diffeomorphism of ${\mathbb{C}}{\mathbb{P}}^{n}$ with more than $n+1$ fixed points must have infinitely many periodic orbits. (Hence, the name “HZ-conjecture”.) Here, by Arnold’s conjecture, $n+1$ is the minimal possible number of fixed points, [Fl89b, Fo, FW]. For instance, $n+1=2$ for ${\mathbb{C}}{\mathbb{P}}^{1}=S^{2}$. The HZ-conjecture was proved in [Sh22] under a minor additional non-degeneracy type condition but for a broader class of manifolds than complex projective spaces; see also, e.g., [Al22a, Al22b, Al23, ABF, AL, BX, Ba15, Ba17, Ba18, ÇGG22, GG14, GG16, Gü13, Su21] for other relevant results, conjectures and more context.
@@
From a slightly different perspective, the $n$-or-$\infty$ conjecture is a question about Reeb pseudo-rotations. In the context of symplectic dynamics, the two competing definitions of a pseudo-rotation are that this is a Hamiltonian system (a Hamiltonian diffeomorphism or a Reeb flow) with finitely many periodic orbits or, alternatively, with the minimal possible number of periodic orbits (perhaps, counted with multiplicity); see [GG18, Sh22, ÇG2M23]. Hypothetically, the two definitions are equivalent and Theorem B establishes this fact for a certain class of Reeb flows, just as [Sh22] did for Hamiltonian diffeomorphisms of ${\mathbb{C}}{\mathbb{P}}^{n}$ and some other manifolds; see also [AL, BX, ÇGG22]. (Interestingly, this is no longer true for symplectomorphisms, [ABF].) Overall, perhaps expectedly so, there is a considerable similarity between the properties of Hamiltonian and Reeb pseudo-rotations; see, e.g., [AK, LRS] vs. [Ka] or [GG18] vs. [ÇG2M23].
@@
Finally, the case of dimension $32n-1=$ 3 deserves a separate discussion. According to the so called 2-or-infinity conjecture, a Reeb flow on a closed 3-manifold has either exactly two or infinitely many prime closed orbits. This conjecture is almost proved in general, and lens spaces are the only manifolds admitting Reeb pseudo-rotations; see [CGHu, CGHP, CGH2L, CDR]. Furthermore, the conjecture has been completely proved for $S^{3}$. (See also [GH2M] for a different proof of (C-M) for $S^{3}$.) Hence, both (C-M) and (C-HZ) parts of our $n$-or-$\infty$ conjecture hold in this case. However, one cannot expect this clear-cut dichotomy to hold in higher dimensions, even for tight or fillable contact manifolds. Indeed, while it is reasonable to expect that a majority of such contact manifolds do not admit Reeb pseudo-rotations in all dimensions just as in dimension three, it is not clear how to rigorously state this conjecture and a description of manifolds that do seems beyond our reach. Moreover, the lower bound on the number of orbits clearly depends on the manifold when $n>2$.
@@
We use Theorem C to show that every homologically visible (over ${\mathbb{Q}}$) closed orbit has one-dimensional equivariant local symplectic homology. This allows to assign such a closed orbit a degree and treat the orbit as if it were non-degenerate. We show that the degrees of such orbits form the increasing sequence $n+1,n+3,\ldots$ and the degrees of bars are $n,n+2,\ldots$. All ${\mathbb{Q}}$-visible closed orbits have distinct period (action) and the same ratio of the action and the mean index. (See Corollary D and Theorem E.) In other words, the (non-equivariant) symplectic homology persistence module and the orbit degrees behave as they would for an irrational ellipsoid.
@@ 1.2. Main results
When $x$ is non-degenerate, we call it non-alternating if the number, counted with algebraic multiplicity, of real negative eigenvalues in $(-1,0)$ of its Poincaré return map is even. Otherwise, we call $x$ alternating. A non-degenerate even iterate of an alternating closed orbit is called bad; all other non-degenerate closed orbits are called good. When $M$ is centrally symmetric, we call $x$ symmetric if $-x(t)=x(t+T/2)$, where $T$ is the period of $x$, i.e., $x$ and $-x$ are the same orbit up to the choice of initial condition.
@@
This theorem is proved in Section 4.1. The key new point here is, of course, (M1), establishing Ekeland’s conjecture and improving previously known multiplicity lower bounds by roughly a factor of two. As we have already mentioned, the existence of at least $n$ non-alternating orbits for non-degenerate dynamically convex Reeb flows is well-known; see [GK, LZ]. The new part in (M2), essential for the proof of Theorem B, is that in this case the number of non-alternating closed prime orbits is exactly $n$ and the remaining prime orbits, if they exist, must all be alternating.
@@ 2.1. Conventions and notation
Recall also that a non-degenerate closed orbit $x$ is said to be non-alternating if the number (counted with algebraic multiplicity) of eigenvalues in $(-1,0)$ of the return map is even. Otherwise, we call $x$ alternating. When $x$ is strongly non-degenerate, these conditions are equivalent to that all indices $\mu\big{(}x^{k}\big{)}$, $k\in{\mathbb{N}}$, have the same parity or that the parity alternates. (A word is due on terminology: Non-alternating orbits are sometimes referred to as negative hyperbolic, but we find this term misleading when $\dim M>3$. The frequently used terms “even/odd” are more intuitive; however they clash with the parity of an orbit.) While formally speaking these definitions require no conditions on $x$, we will apply them only when $x$ is prime and strongly non-degenerate. The notion of (non-)alternating closed orbit is not particularly useful when $x^{k}$ is allowed to be degenerate; see Example 2.13. The parity of a closed non-degenerate orbit $x$ is simply the parity of $\mu(x)$. A non-degenerate closed orbit $x$ is said to be bad if it is an even iterate of an alternating prime orbit, and good otherwise.
@@
This condition is obviously satisfied in two cases we are interested in here: when $\alpha$ is non-degenerate and when the Reeb flow has only finitely many prime closed orbits. The role of this condition is rather expository than conceptual, and most of the discussion in the next two sections carries over to the general case with only technical changes. For instance, the symplectic homology is still a persistence module (see Section 2.2) when ${\mathcal{S}}$ is only closed and nowhere dense and $V^{s}$ might be infinite-dimensional, once (PM2) in the definition of a persistence module is suitably modified; [ÇG2M24].
@@ 2.3. Filtered symplectic homology
The construction of symplectic homology carries over to ${\mathbb{Z}}$ or any other coefficient ring, and hence we have the universal coefficient theorem. In particular, for a fixed $t$, in every degree $m$ we have
@@
Recall in this connection that $\operatorname{SH}^{\infty}(W;{\mathbb{F}})$ vanishes, for instance, whenever $W$ is displaceable in $\widehat{W}$. For example, $\operatorname{SH}(W)^{\infty}=0$ when $W$ is a star-shaped domain in $\widehat{W}={\mathbb{R}}^{2n}$. In this case, vanishing of $\operatorname{SH}^{\infty}(W)$ is established in [Vi99]. The general case is proved in [CFO] via Rabinowitz–Floer homology and direct proofs are given in [Su16] and [GS]. Strictly speaking, Theorem 2.3 is proved in [GS] for ${\mathbb{F}}={\mathbb{Q}}$, but the argument carries over word-for-word to any unital ring.
@@
It is worth keeping in mind that this description does not carry over to a field of positive characteristic $p\geq 2$. The reason is that $x^{p}$ can contribute infinitely many generators to the complex over ${\mathbb{F}}_{p}$. For instance, if $x^{p}$ is non-degenerate and the only orbit with action in $I$ and $x$ is prime, $\operatorname{CH}^{I}(W;{\mathbb{F}}_{p})$ is isomorphic to the homology of the infinite-dimensional lens space $S^{2\infty-1}/{\mathbb{Z}}_{p}$ over ${\mathbb{F}}_{p}$.
@@ 2.4.2. Behavior under iterations
The notion of an (non-)alternating closed orbit is not particularly useful in the setting where $x$ is not strongly non-degenerate and the return map $\varphi$ has roots of unity of even degree among its eigenvalues. The reason is that in this case the behavior of the local symplectic homology is not determined by the eigenvalues and depends on higher order terms. For instance, in the spirit of Example 2.12, let $R_{\pi}\colon{\mathbb{R}}^{2}\to{\mathbb{R}}^{2}$ be the counterclockwise rotation by $\pi$, i.e., central symmetry, viewed as an element of $\operatorname{\widetilde{Sp}}(2)$. Consider the Hamiltonians $H_{0}=p^{4}-q^{4}$ and $H_{1}=(p^{2}+q^{2})^{2}$ on ${\mathbb{R}}^{2}$, where $(p,q)$ are the Darboux coordinates. The flows of $H_{0}$ and $H_{1}$ commute with $R_{\pi}$. Let $\varphi_{i}$, for $i=0,1$, be obtained by composing $R_{\pi}$ with the time-one flow of $H_{i}$ and $x_{i}$ be a closed Reeb orbit with the return map $\varphi_{i}$. Clearly, $\operatorname{CH}(x_{i};{\mathbb{Q}})=\operatorname{HF}(\varphi_{i})={\mathbb{Q}}$ is supported in degree one. However, one can show that $\operatorname{CH}\big{(}x_{0}^{2};{\mathbb{Q}}\big{)}=0$ and $\operatorname{HF}\big{(}\varphi_{0}^{2};{\mathbb{Q}}\big{)}={\mathbb{Q}}$ (supported in degree two). At the same time, $\operatorname{CH}\big{(}x_{1}^{2};{\mathbb{Q}}\big{)}=\operatorname{HF}\big{(}\varphi_{1}^{2};{\mathbb{Q}}\big{)}={\mathbb{Q}}$ (supported in degree three).
@@ 2.4.3. From local to filtered homology
where, when $0\in I$, we treat $W$ as a collection of “orbits” with action 0 as described above. This spectral sequence is simply associated with the action filtration; see, e.g., [GG20, Sec. 2].
@@ 2.5. The beginning and end maps
For a fixed field ${\mathbb{F}}$, denote by ${\mathcal{B}}$ or ${\mathcal{B}}({\mathbb{F}})$ the collection of bars of the graded persistence module $\operatorname{SH}(W;{\mathbb{F}})$. Assume that $\operatorname{SH}^{\infty}(W;{\mathbb{F}})=0$. Then every bar in ${\mathcal{B}}$ is finite by Theorem 2.3. Furthermore, let ${\mathcal{P}}={\mathcal{P}}({\mathbb{F}})$ be the set of ${\mathbb{F}}$-visible closed Reeb orbits. As in the previous section, let us also formally add to ${\mathcal{P}}$ the set of the “beginning” of the bars of the form $I=(0,a]$ for some $a>0$, denoting the resulting set by ${\mathcal{P}}^{\prime}$. These additional elements $x$ of ${\mathcal{P}}^{\prime}$ are simply in one-to-one correspondence with such bars. By definition, each of such “orbits” $x$ has symplectic homology equal to ${\mathbb{F}}$ and concentrated in $\deg(I)$, and zero action. (We also set $\mu_{-}(x)=\mu_{+}(x):=\deg(x)$ for such $x$.) When $W$ is star-shaped, we are adding just one extra element corresponding to the domain $W$, which we denote by $[W]$. This element has action zero and its supported in degree $n$.
@@ 3.1.3. Conley–Zehnder index: closed orbits
Next, let $x\colon S_{T}^{1}:={\mathbb{R}}/T{\mathbb{Z}}\to M$ be a closed orbit of the Reeb flow, which we do not require to be prime. The pull-back ${\mathfrak{s}}_{M}\circ x$ of ${\mathfrak{s}}_{M}$ by $x$ is a non-vanishing section of the pull-back $x^{*}L_{M}$ over $S_{T}^{1}$ and $x^{*}L_{M}$ is in turn the top complex wedge of $x^{*}\xi$. Pick any Kähler trivialization of $x^{*}\xi$ which, up to homotopy, gives rise to the trivialization ${\mathfrak{s}}_{M}\circ x$ of $x^{*}L$, i.e., such that its top complex wedge is ${\mathfrak{s}}_{M}\circ x$. (Such a trivialization is unique up to homotopy.) Then we use this trivialization to turn the linearized Reeb flow $D\varphi^{t}$, $t\in[0,T]$, along $x$ into a path $\Phi(t)\in\operatorname{Sp}(2m)$ with $m=n-1$ starting at the identity, and hence an element $\Phi\in\operatorname{\widetilde{Sp}}(2m)$. By definition, $\mu_{\pm}(x):=\mu_{\pm}(\Phi)$ and $\operatorname{\hat{\mu}}(x):=\operatorname{\hat{\mu}}(\Phi)$. In particular, when $x$ is non-degenerate, $\mu(x):=\mu(\Phi)$. Furthermore, $\operatorname{\hat{\mu}}$ is homogenenous, i.e.,